# Collective Intelligence Outperforms Individual Talent: A Case Study in League of Legends

### v1

@@ Abstract
Gaming environments are popular testbeds for studying human interactions and behaviors in complex artificial intelligence systems. Particularly, in multiplayer online battle arena (MOBA) games, individuals collaborate in virtual environments of high realism that involves real-time strategic decision-making and trade-offs on resource management, information collection and sharing, team synergy and collective dynamics. This paper explores whether collective intelligence, emerging from cooperative behaviours exhibited by a group of individuals, who are not necessarily skillful but effectively engage in collaborative problem-solving tasks, exceeds individual intelligence observed within skillful individuals. This is shown via a case study in League of Legends, using machine learning algorithms and statistical methods applied to large-scale data collected for the same purpose. By modelling systematically game-specific metrics but also new game-agnostic topological and graph spectra measures of cooperative interactions, we demonstrate compelling insights about the superior performance of collective intelligence.
@@ 1 Introduction
The recent advancements in computing technology have evolved the way individuals socialize and interact; online gaming platforms being one of the most prominent mediums for social interaction and cooperation. Zhong reports that in China there are 147 million people playing online games, thus representing new forms of community, social interaction, and collaboration [1]. These online gaming environments are popular testbeds for studying human interactions and behaviours in complex artificial intelligence systems. The simulation of diverse scenarios ranging from strategic decision-making to adaptive behaviours within online gaming environments allows the testing of complex real-world mechanisms under controlled low-cost conditions. Furthermore, the virtual worlds of such online games are populated with autonomous artificial agents and other human players, thus offering opportunities for interesting interactions that guide players and enhance social dynamics [2], resulting in a collective intelligence derived from the synergistic cooperation of humans and computing agents. Collective intelligence is the general ability of a particular group to perform well across a wide range of different tasks [3]. In the field of computer science, it involves the study of how groups of one person and one computer interact (human-computer interaction) as well as how larger groups of people interact with computing technologies to cooperate and make intelligent decisions, for instance, in the development of collaborative software such as Wikipedia [4].
@@
Collective intelligence is prominently demonstrated in many online gaming environments among which multiplayer online battle arena (MOBA) games serve as a notable example. MOBA games are a subgenre of real-time strategy games in which two teams, each consisting of 5 individual players, compete against each other with each player controlling a single character (called a champion), while strategy revolves around individual character development and cooperative team play in combat [5]. Riot Games’ League of Legends is the most popular MOBA game played worldwide with over 150 million monthly active players [6]. League of Legends features 169 uniquely playable champions with their own set of skills, that relies on an individual player’s micro-management ability, paired with a strong understanding of champion mechanics along with the collective effort of a team to think strategically and showcase team-work abilities. The game is structured around a map called Summoner’s Rift, and the primary objective of the game is to destroy the opponent team’s base, a tower structure called the Nexus, while the game progresses with champion development via earning gold, gaining experience, crafting items and scoring kills.
The collaborative dynamics observed among randomly assigned players within a team, each controlling a distinct champion with unique sets of abilities and working towards the shared objective of winning the game, potentially exemplifies the presence of collective intelligence in MOBA games. In fact, exchange of in-games chat messages, signals and assists to secure game objectives elicits a quantitative measure of collective intelligence. Leavitt et al. show that non-verbal communication in the form of in-game pings improves team strategy and performance [7]. On the other hand, individual players can excel in executing their champion mechanics and abilities, thus displaying high levels of individual intelligence. Using in-game data to model various metrics of individual and collective intelligence, this paper aims to outline the importance of such individual as well as team-based collective intelligence metrics in influencing the outcome of a game. That is, it tests the hypothesis H1: collective intelligence outperforms individual intelligence of the players in winning a game. Additionally, Riot Games maintains a two-tiered in-game ranking system (called tier and division) that establishes a hierarchy of players based on their recent in-game performances. Using such information, this paper also tests another hypothesis, that is, H2: Highly-skilled players exhibit higher levels of collective intelligence compared to less-skilled players. The main contributions of this work are – (i) a framework to evaluate individual and collective intelligence of players in League of Legends using in-game metrics and machine learning algorithms, and (ii) the availability of dataset and source code that can be used to reproduce the results.
@@ 3 Methods
To evaluate the aforementioned hypotheses within the League of Legends gameplay environment, different statistical methods and machine learning algorithms are implemented. Figure 1 depicts the overall framework of the methods used in this study. Individual in-game performance metrics are extracted from the experimental dataset while team-based collective intelligence metrics are computed from their corresponding directed graph networks that outline the flow of resources between players in a team. Exploratory factor analysis is applied to both individual and team-level data, revealing latent performance factors. Such latent factors are input as features to predict the winning outcome of a game using different classification algorithms in order to understand the nature of such intelligent behaviours. Comparison of the classification results based on the individual and collective performance factors determine whether or not to accept the hypothesis H1. Furthermore, k-means clustering is used to profile players and teams into different clusters. Crosstabs analysis of the winning rate of different teams having different player clusters reinforces whether or not to accept the hypothesis H1. Finally, statistical hypothesis testing on H2 is employed to check for the significance in the difference of collective intelligence levels of teams (and players) across different ranks. Details for each step are further outlined as follows–
Figure 1: Overall architectural framework for the evaluation of collective intelligence hypotheses within League of Legends gameplay environments.
@@ 3.1 Experimental Dataset
League of Legends game data played on official servers is publicly accessible via a set of RESTful APIs provided by Riot Games [27]. In order to fetch game data for competitive matches played by a given player, a summonerName which uniquely identifies the player, is required. A preliminary list consisting of 10 random summonerNames corresponding to the 10 different ranking tiers, is used as initial seed data. For each of these seed summonerNames, the API offered by Riot Games is used to retrieve – (a) a list of $N=100$ recent matches played by the player, and for each of these matches, (b) detailed match data that include game outcome and individual match performance of the players, (c) a list of game events that occurred at periodic time intervals within the match and, (d) public rankings of the associated players. The API fetches not just the match performance of the seeded summonerName player, but also retrieves details for the other 9 players relevant in a given match. An additional set of data i.e., the most recent $N=10$ games for all the relevant players identified from the initial fetch, is collected, thus removing any initial selection bias. Using the process described above, a total of 103K matches are fetched. However, due to different modes of gameplay offered in the game, the collected dataset does not exclusively consist of ranked competitive matches only. Furthermore, two different players in this dataset may have played in the same game, or there can be games where a player was disconnected (due to Internet connection issues, or abandoning the game) from the server resulting in less than 10 players, or the game was terminated early via a remake or a surrender. To remedy such issues, the dataset was filtered to remove duplicate entries, retrieving only those games in which all 10 players were present throughout the game in a competitive ranked match mode that was not terminated early (i.e., game duration is not less than 23 minutes). Thus the experimental dataset consisted a total of 31,055 ranked matches with 164,571 unique players (see Appendix for details on data-fetch process). For each game, in-game performance metrics of all players were extracted and directed graph networks were generated to formulate team-based graph metrics.
@@ 3.2 In-game Performance Metrics
The dataset collected from the API endpoints, particularly the match detail endpoint, consists of fields that reflect the respective in-game performance of all players in a match in terms of metrics such as – experience gained, gold earned, kills score, assists given, champion deaths, vision score, elite monsters kills, towers destroyed, minions killed, etc. along with the overall match outcome. Feature scaling was performed to resolve the discrepancies in the range of values for the in-game metrics. Additionally, per-minute metrics for gold earned, assists given, vision score, experience gained and minions killed were also computed to scale and normalize the range of features in the dataset. Additionally, champions played by individual players feature their own set of unique skillshots, spells and items used during the gameplay which can impact the outcome of a game. However, incorporating such features only adds complexity without providing high-level insights on the nature of intelligent behaviors. Thus champion skillshots, spells and item builds are not accounted as in-game performance metrics for simplicity. Details on the gameplay and in-game metrics is further explained in Appendix.
@@ 3.3 Graph Metrics
Graph metrics characterize the flow and centrality of resources between individuals in a team. Such graph metrics signify how teams and individuals interact, strategize and cooperate throughout the game to secure different in-game objectives and resources. To explore the interaction among players, we utilize the match timeline data to fetch game events during which players in a team assisted one another to secure objectives and resources such as opponent kills, enemy tower destruction, elite monster kills as well as unrecorded assists for positional map-pressure resulting in a kill. Positional map-pressure refers to the intelligent positioning of a player on the map to prevent an opponent player from escaping, forcing its kill without actively participating in it (see Appendix).
@@
Player In-Degree. This metrics signifies the weighted sum of assists received by a player from all other players in the team i.e. the in-degree of player $i$ is: $C_{ID}(i)=\sum_{j=1}^{N}w_{ji}$
Player Out-Degree. This metric signifies the weighted sum of assists given by a player to all other players in the team i.e. the out-degree of player $i$ is: $C_{OD}(i)=\sum_{j=1}^{N}w_{ij}$
Team In-Degree Centrality. This metric signifies the extent to which assists are centralized towards any player in a given team i.e.: $C_{I}=\frac{\sum_{i=1}^{N}(C_{ID}^{max}-C_{ID}(i))}{(N-1)A}$
Team Out-Degree Centrality. This metric signifies the extent to which assists are given out by any player in a given team i.e.: $C_{O}=\frac{\sum_{i=1}^{N}(C_{OD}^{max}-C_{OD}(i))}{(N-1)A}$
Effective Graph Resistance. This metric measures the robustness of a graph network [28]. A higher value of this metric signifies the lack of cooperation among nodes in the graph network. This metric is computed by extracting eigenvalues ($\mu$) from the corresponding Laplacian matrix of the graph network for each team i.e.: $EGR=N\sum_{i=1}^{N-1}\frac{1}{\mu_{i}}$
@@ 3.4 Individual Metrics vs Collective Metrics
An individual’s performance in a game is accounted by the in-game performance metrics of the individual player such as gold earned, champion experience, kills, etc. Correlation analysis and variance inflation factor tests revealed multi-collinearity between in-game metrics which were thus removed from the feature set. Additionally, the graph metrics i.e. player in-degree and player out-degree which signify the overall receipt and provision of all assists for a single individual player is also accounted towards individual metrics. Thus, the individual metrics include – vision score per minute, gold earned per minute, champion experience per minute, player in-degree and player out-degree. The collective metrics are composed of average values of in-game performance metrics of the players in a team along with the graph metrics that signify the overall flow of assistance within the team. Thus, the collective metrics consist of average team gold earned per minute, average team vision score per minute, average team champion experience per minute, team in-degree centrality, team out-degree centrality and effective graph resistance.
@@ 3.6 Machine Learning Algorithms
Classification: The winning outcome of a game in most cases is determined either by the exceptional performances of an individual player or by the collective efforts of a team. The explanatory factors that suggest the nature of intelligent behaviours in individuals and teams are then input as features to predict the winning outcome of the game using different classification algorithms i.e., binary logistic regression, decision tree classifier, random forest classifier, and XGBoost classifier. Classification results are evaluated in terms of accuracy, f1-score and area under the curve metrics. To accept or reject H1, the evaluation scores by the collective performance factors are compared against individual performance factors.
Clustering: Further analysis involves profiling individual players and teams based on the explanatory factors into three distinct clusters for each. K-means clustering algorithm is used to categorize individual players and teams into their respective clusters. Based on the average values of the performance factors within each cluster, we manually add labels to these clusters. The frequency of wins (or winning rate) within such distinct clusters of players and teams are then identified. Cross tabulation analysis enables how often a team cluster with a majority of particular type of players win the game. Comparison of the winning rates of such player and teams clusters via cross-tabulation analysis provides further rigour to accept or reject the hypothesis H1.
@@ 4.1 Individual Acquisition but Collective Cooperation
Exploratory factor analysis on the individual and team collective metrics reveal latent factors which relate to acquiring vs sharing of resources in individuals and cooperative vs non-cooperative behaviors in teams. The optimal number of factors for both individual and collective team performance metrics were identified to be $N=2$ via Scree analysis. Based on the factor loading values of the latent factors, such factors could be interpreted as either acquiring or sharing of resources in individuals, and as either cooperative or non-cooperative factors in teams. Table 1 outlines the factor loading values of underlying components for both individual and teams. Upon training the match dataset represented by such latent factors at both the individual and team-level to predict the winning outcome using different classification algorithms, the feature importance values for such latent factors present an interesting dynamic between individual and collective behaviour i.e., individuals attempt to win games via acquisition of resources for a player whereas teams attempt to win games via cooperative, altruistic behaviors. Figure 3 presents the feature importance values of the latent factors at the individual and collective team-level across different sets of classification algorithms.
Table 1: Factor loading values of the latent factors for individual and collective team metrics. Individual performance metrics shaded in gray benefit another player more than themselves, while non-shaded metrics are related to acquiring of resources for a player. The factor loading values for Factor 1 shows high-positive correlation with acquiring and low-negative correlation with sharing of resources, and vice-versa for Factor 2. Thus, Factor 1 and Factor 2 are distinguished as acquiring and sharing factor respectively. Likewise, collective team metrics shaded in gray relate to the resistance in the flow (or centralization) of resources and assistance within the team. The factor loading values for Factor 1 shows low positive correlation values in centralization of resources and high negative correlation for resource acquisition (average gold and experience), while Factor 2 shows higher positive correlation with metrics that signify centralization. Thus Factor 1 and Factor 2 is distinguished as cooperative and non-cooperative factor respectively.

### v2

@@ Abstract
Gaming environments are popular testbeds for studying human interactions and behaviors in complex artificial intelligence systems. Particularly, in multiplayer online battle arena (MOBA) games, individuals collaborate in virtual environments of high realism that involves real-time strategic decision-making and trade-offs on resource management, information collection and sharing, team synergy and collective dynamics. This paper explores whether collective intelligence, emerging from cooperative behaviours exhibited by a group of individuals, who are not necessarily skillful but effectively engage in collaborative problem-solving tasks, exceeds individual intelligence observed within skillful individuals. This is shown via a case study in League of Legends, using machine learning algorithms and statistical methods applied to large-scale data collected for the same purpose. By modeling and visualizing systematically game-specific metrics but also new game-agnostic topological and graph spectra measures of cooperative interactions, we demonstrate compelling insights about the superior performance of collective intelligence.
@@ 1 Introduction
The recent advancements in computing technology have evolved the way individuals socialize and interact; online gaming platforms being one of the most prominent mediums for social interaction and cooperation. Zhong reports that in China there are 147 million people playing online games, thus representing new forms of community, social interaction, and collaboration [1]. These online gaming environments are popular testbeds for studying human interactions and behaviors in complex artificial environments. The simulation of diverse scenarios ranging from strategic decision making to adaptive behaviors within online gaming environments allows the testing of complex real-world mechanisms under controlled low-cost conditions. Furthermore, the virtual worlds of such online games are populated with autonomous artificial agents and other human players, thus offering opportunities for interesting interactions that guide players and enhance social dynamics [2], resulting in a collective intelligence derived from the synergistic cooperation of humans and computing agents. Collective intelligence is the general ability of a particular group to perform well across a wide range of different tasks [3]. Collective intelligence also involves the study of how groups of one person and one computer interact (human-computer interaction) as well as how larger groups of people interact with computing technologies to cooperate and make intelligent decisions, for instance, in the development of collaborative software such as Wikipedia [4].
@@
MOBA games are a subgenre of real-time strategy games in which two teams, each consisting of 5 individual players, compete against each other with each player controlling a single character (called a champion), while strategy revolves around individual character development and cooperative team play in combat [5]. Riot Games’ League of Legends is the most popular MOBA game played worldwide with over 150 million monthly active players [6]. League of Legends features 171 uniquely playable champions with their own set of skills, that relies on an individual player’s micro-management ability, paired with a strong understanding of champion mechanics along with the collective effort of a team to think strategically and showcase team-work abilities. The game is structured around a map called Summoner’s Rift, and the primary objective of the game is to destroy the opponent team’s base, a tower structure called the Nexus, while the game progresses with champion development via earning gold, gaining experience, crafting items and scoring kills (refer Figure 1).  [^1] Figure 1A is sourced from Marcal Mora Cantallops [7], while Figure 1B and Figure 1C are re-worked from in-game interfaces
Figure 1: (A) The three lanes, 2 jungles and their subdivisions. Small red and blue circles denote tower structures for the respective teams. The darker, inner-circle segment in the top-right and bottom-left corners denote each teams’ Nexus; (B) An instance of gameplay on Summoner’s Rift where an opponent kill is scored due to positional map-pressure. Blue team player A attempts to kill red team player Y, who tries to flee. Blue player B blocks Y’s escape, forcing engagement and leading to A’s kill. Although B does not actively assist, B’s positioning creates map pressure resulting in Y’s death. B is awarded a map-pressure assist if it is positioned within a specified radius during the kill. (C) Details of an in-game champion;
The collaborative dynamics among randomly assigned players in a team, each controlling a unique champion and working toward a shared win objective, exemplify collective intelligence in MOBA games. In-game chat messages, signals, and assists used to secure objectives offer measurable indicators of collective intelligence. Leavitt et al. show that nonverbal communication like in-game pings improves team strategy and performance [8]. Meanwhile, individual players may excel in champion mechanics, reflecting high individual intelligence. Extending existing work [9], this paper uses in-game data to model and visualize metrics of both individual and collective intelligence, testing the hypothesis H1: collective intelligence outperforms individual intelligence of the players in winning games. These players in such competitive games are categorized into a two-tiered ranking system of tier and division; based on recent performance, serving as proxies for individual skill levels that influence outcomes. This motivates H2: Highly-skilled players exhibit higher levels of collective intelligence compared to less-skilled players. Although player perceptions of in-game incentives affect skill development [10], such mechanisms are beyond the scope of this study. The main contributions of this work in evaluating individual and collective intelligence in League of Legends are: (i) the first use of effective graph resistance from graph spectra as a proxy for collective intelligence, previously used for network robustness [11]; (ii) new visual communication of collective intelligence by combining graph and complex network approaches with chord diagrams; and (iii) new insights into how collective intelligence appears in League of Legends and how it compares with individual player intelligence.
@@ 3 Methods
To evaluate the aforementioned hypotheses within the League of Legends gameplay environment, different statistical methods and machine learning algorithms are implemented. Figure 2 depicts the overall framework of the methods used in this study. Individual in-game performance metrics are extracted from the experimental dataset while team-based collective intelligence metrics are computed from their corresponding directed graph networks that outline the flow of resources between players in a team. Exploratory factor analysis is applied to both individual and team-level data, revealing latent performance factors. Such latent factors are input as features to predict the winning outcome of games using different classification algorithms in order to understand the nature of such intelligent behaviours. Comparison of the classification results based on the individual and collective performance factors determine whether or not to accept the hypothesis H1. Furthermore, k-means clustering is used to profile players and teams into different clusters. Crosstabs analysis of the winning rate of different teams having different player clusters reinforces whether or not to accept the hypothesis H1. Finally, statistical hypothesis testing on H2 is employed to check for the significance in the difference of collective intelligence levels of teams (and players) across different ranks in the case of team wins.
Figure 2: Overall architectural framework for the evaluation of collective intelligence hypotheses within League of Legends gameplay environments. The architecture comprises two main components: (i) computation of individual and team metrics, and (ii) analysis of winning outcomes using exploratory data analysis with classification and clustering methods on such individual and collective team-level data.
@@ 3.1 Experimental Dataset
League of Legends game data played on official servers is publicly accessible via a set of RESTful APIs provided by Riot Games [31]. In order to fetch game data for competitive matches played by a given player, a summonerName which uniquely identifies the player, is required. A preliminary list consisting of 10 random summonerNames corresponding to the 10 different ranking tiers, is used as initial seed data. For each of these seed summonerNames, the API offered by Riot Games is used to retrieve – (a) a list of $N=100$ recent matches played by the player, and for each of these matches, (b) detailed match data that include game outcome and individual match performance of the players, (c) a list of game events that occurred at periodic time intervals within the match and, (d) public rankings of the associated players. The API fetches not just the match performance of the seeded summonerName player, but also retrieves details for the other 9 players relevant in a given match. An additional set of data i.e., the most recent $N=10$ games for all the relevant players identified from the initial fetch, is collected, thus removing any initial selection bias. Using the process described above, a total of 103K matches are fetched. However, due to different modes of gameplay offered in the game, the collected dataset does not exclusively consist of ranked competitive matches only. Furthermore, two different players in this dataset may have played in the same game, or there can be games where a player was disconnected (due to Internet connection issues, or abandoning the game) from the server resulting in less than 10 players, or the game was terminated early via a remake or a surrender. To remedy such issues, the dataset was filtered to remove duplicate entries, retrieving only those games in which all 10 players were present throughout the game in a competitive ranked match mode that was not terminated early (i.e., game duration is not less than 23 minutes preventing any early surrender and ensuring both teams are competing to win towards the latter stages of the game regardless of the early dominance by a team). Thus the experimental dataset consisted a total of 31,055 ranked matches with 164,571 unique players. For each game, in-game performance metrics of all players were extracted and directed graph networks were generated to formulate team-based graph metrics.
@@ 3.2 In-game Performance Metrics
The dataset collected from the API endpoints, particularly the match detail endpoint, consists of fields that reflect the respective in-game performance of all players in a match in terms of metrics such as – experience gained, gold earned, kills score, assists given, champion deaths, vision score, elite monsters kills, towers destroyed, minions killed, etc. along with the overall match outcome. Feature scaling was performed to resolve the discrepancies in the range of values for the in-game metrics. Additionally, per-minute metrics for gold earned, assists given, vision score, experience gained and minions killed were also computed to scale and normalize the range of features in the dataset. Additionally, champions played by individual players feature their own set of unique skillshots, spells and items crafted during the gameplay which can impact the outcome of a game. However, incorporating such features only adds complexity without providing high-level insights on the nature of intelligent behaviors. Thus champion skillshots, spells and item builds are not accounted as in-game performance metrics for simplicity.
@@ 3.3 Graph Metrics
Graph metrics characterize the flow and centrality of resources between individuals in a team. Such graph metrics signify how teams and individuals interact, strategize and cooperate throughout the game to secure different in-game objectives and resources. To explore the interaction among players, we utilize the match timeline data to fetch game events during which players in a team assisted one another to secure objectives and resources such as opponent kills, enemy tower destruction, elite monster kills as well as unrecorded assists for positional map-pressure resulting in a kill. Positional map-pressure refers to the intelligent positioning of a player on the map to prevent an opponent player from escaping, forcing its kill without actively participating in it.
@@
1. Player In-Degree: This metrics signifies the weighted sum of assists received by a player from all other players in the team i.e. the in-degree of player $i$ is: $C_{ID}(i)=\sum_{j=1}^{N}w_{ji}$.
2. Player Out-Degree: This metric signifies the weighted sum of assists given by a player to all other players in the team i.e. the out-degree of player $i$ is: $C_{OD}(i)=\sum_{j=1}^{N}w_{ij}$.
3. Team In-Degree Centrality: This metric signifies the extent to which assists are centralized towards any player in a given team i.e.: $C_{I}=\frac{\sum_{i=1}^{N}(C_{ID}^{max}-C_{ID}(i))}{(N-1)A}$.
4. Team Out-Degree Centrality: This metric signifies the extent to which assists are given out by any player in a given team i.e.: $C_{O}=\frac{\sum_{i=1}^{N}(C_{OD}^{max}-C_{OD}(i))}{(N-1)A}$.
5. Effective Graph Resistance (EGR): This metric measures the robustness of a graph network [11]. A higher value of this metric signifies the lack of cooperation among nodes in the graph network. This metric is computed by extracting eigenvalues ($\mu$) from the corresponding Laplacian matrix of the graph network for each team i.e.: $EGR=N\sum_{i=1}^{N-1}\frac{1}{\mu_{i}}$.
The graph metrics – player in-degree, player out-degree, team in-degree centrality, team out-degree centrality have been used in previous work [28], but our approach models all forms of assistance i.e. assistance on kills including towers taken, elite-monsters and positional map-pressure
@@ 3.4 Individual Metrics vs Collective Metrics
An individual’s performance in a game is accounted by the in-game performance metrics of the individual player such as gold earned, champion experience, kills, etc. Correlation analysis and variance inflation factor tests revealed multi-collinearity between in-game metrics which were thus removed from the feature set. Additionally, the graph metrics i.e. player in-degree and player out-degree which signify the overall receipt and provision of all assists for a single individual player is also accounted towards individual metrics. The collective metrics are composed of average values of in-game performance metrics of the players in a team along with the graph metrics that signify the overall flow of assistance within the team. Table 1 outlines the list of individual metrics and collective metrics used in the study.
Table 1: Performance metrics of individuals and teams
Individual Metrics Collective Metrics
Gold Earned Per Min. Avg. Gold Earned Per Min.
Experience Per Min. Avg. Experience Per Min.
Vision Score Per Min. Avg. Vision Score Per Min.
Player In-degree Team In-degree Centrality
Player Out-degree Team Out-degree Centrality
Effective Graph Resistance
@@ 3.6 Machine Learning Algorithms
Classification: A game’s winning outcome is often driven by either exceptional individual performance or collective team effort. Explanatory factors reflecting intelligent behaviors in individuals and teams are used as features to predict the game’s outcome through classification algorithms—binary logistic regression, decision tree, random forest, and XGBoost. Results are evaluated using accuracy, f1-score, and area under the curve. To assess H1, evaluation scores from collective performance factors are compared to those from individual ones.
Clustering: Additional analysis profiles players and teams into three clusters each, based on explanatory factors, using k-means clustering. Clusters are then manually labeled based on the average values of performance factors in each cluster. Win frequencies within these clusters are identified. Cross-tabulation shows how often teams with a majority of a certain player type win. Comparing these win rates across player and team clusters supports testing H1.
@@ 4.1 Individual Acquisition but Collective Cooperation
Exploratory factor analysis on individual and team collective metrics reveals latent factors related to acquiring versus sharing in individuals, and cooperative versus noncooperative behaviors in teams. Scree analysis identifies the optimal number of factors as $N=2$ for both levels. Based on factor loading values, individual factors reflect either resource acquisition or sharing, while team factors represent either cooperation or noncooperation. Table 2 shows these loading values. Using these latent factors, classification models are trained to predict game outcomes. Feature importance scores reveal a dynamic: individuals tend to win through resource acquisition, while teams win through cooperative behaviors. At the individual level, acquisition is more critical [score = 0.68], however at the team level, cooperation and sharing [score = 0.82] are more influential than noncooperative strategies.
Table 2: Factor loading values of latent factors for individual and collective team metrics are shown. For individual metrics, gray-shaded ones benefit others more than the player, while non-shaded ones reflect self-resource acquisition. Factor 1 shows high positive correlation with resource acquisition and low negative with sharing, and vice versa for Factor 2, distinguishing them as acquiring and sharing factors. For collective metrics, gray shading indicates resistance to the flow or centralization of resources and assistance. Factor 1 shows low positive correlation with centralization and high negative with resource acquisition (average gold and experience), while Factor 2 shows higher positive correlation with centralization metrics. Thus, Factor 1 and Factor 2 are identified as cooperative and noncooperative factors respectively.